OpenStudy (anonymous):

how to do this? Find the scalar equation of the plane that is perpendicular to the plane x + 2y + 4 = 0, contains the origin, and whose normal makes an angle of 30 with the z-axis.

OpenStudy (anonymous):

i know that the normal of the required plane is perpendicular to the normal of the given plane, so their dot product must be zero i let <a,b,c> be the normal of the required plane so i got a+2b=0

OpenStudy (anonymous):

-(2/√5) x + (1/√5) y + √3 z =0 this is the answer

OpenStudy (anonymous):

I'm just here for a medal *_* sorry @Data_LG2 but your question is Greek to meh!

OpenStudy (wolfe8):

Hmm I'm having a problem drawing. When I click 'draw', it posts whatever I have typed in. But if you sketch the line, you can imagine how the plane looks like. So I guess you can find the line that's perpendicular to this line and goes through (0,0), and somehow just add a z component? That z component must make an angle to z-axis

OpenStudy (wolfe8):

I would, for the z part, imagine/sketch it on a z-y or z-x plane. I'm sure I'm making no sense.

OpenStudy (anonymous):

i did this: http://sketchtoy.com/58404478 but it's wrong :/

OpenStudy (ikram002p):

this is the plane right x + 2y + 4 = 0 so its normal is <1,2,0>

OpenStudy (anonymous):

yes

OpenStudy (ikram002p):

lets assume the normal of the plane u wanna is <a,b,c>

OpenStudy (anonymous):

okay, i did that.... then?

OpenStudy (ikram002p):

sry i dint know u do it then <a,b,c>.<1,2,0>=0 so a+2b+0c=0 right ?? this is equation 1

OpenStudy (anonymous):

that's fine :) yes..i posted my solution before, but i still don't get the right answer here: http://sketchtoy.com/58404478 and.....?

OpenStudy (ikram002p):

ohk :) u have the angle of 30 with the z-axis. which is DIRECTION ANGLE whith z axis that mean cos 30 = c /||n||

OpenStudy (anonymous):

mhmm

OpenStudy (ikram002p):

||n||=c/cos 30 ||n|| =root a^2+b^2+c^2 c/cos 30=root a^2+b^2+c^2 equation 2 got it ??

OpenStudy (ikram002p):

which part confusing u ? ill explain

OpenStudy (anonymous):

i understand it, but i did not get the right answer, that's why i'm asking how to do it maybe my method is wrong so what will be your next step?

OpenStudy (ikram002p):

so u have 2 equation a+2b=0 c^2=3a^2+3b^2 right ?

OpenStudy (ikram002p):

aasume c humm any num u wanna then solve for a and b

OpenStudy (anonymous):

wait! how did you come up with c^2=3a^2+3b^2 from c/cos 30=root a^2+b^2+c^2??

OpenStudy (anonymous):

aahh got it..

OpenStudy (ikram002p):

cos 30 = root 3 /2 c/ cos 30 = 2c/root 3=root a^2+b^2+c^2 take ^2 to both side 4 c^2 /3 =a^2+b^2+c^2 4 c^2= 3a^2+3b^2+3c^2 c^2= 3a^2+3b^2

OpenStudy (anonymous):

so any number for "c"??

OpenStudy (ikram002p):

any number of c take 2 /root 3 for example it will give u unit normal or take any integer u eant it dosnt mater

OpenStudy (anonymous):

ok, give me a sec, i'll try to solve it

OpenStudy (ikram002p):

becouse of this <a,b,c>.<1,2,0>=0 a+2b+0c=0 so any num of c wont be wrong

OpenStudy (anonymous):

so i assume c=4 i solved for b, and i got 4/√15 and for a i got -8/√15 which means the equation will be: -8/√15 x + 4/√15 y + 4z =0? if i simplified it , it will be -2/√15 x + 1/√15 y + z =0 ALMOST same with the right answer but instead of √15 it should be √5 and the z should have a coefficient of √3 .. the right answer is : -(2/√5) x + (1/√5) y + √3 z =0 ..mm.. should i multiply the equation with something to get it √5 ?

OpenStudy (ikram002p):

humm wait ill check

OpenStudy (ikram002p):

assume c= root 3 it will give u the same answer

OpenStudy (ikram002p):

c=√3 3a^2+3b^2=3 a^2+b^2=1 a=-2b (-2b)^2+b^2=1 b= (1/√5) a=-(2/√5)

OpenStudy (ikram002p):

but yess if u simplify u answer ull get the same equation

OpenStudy (anonymous):

ooh.. i think it's the same thing!! for my answer before, i just need to multiply 1/√5 to get the same equation and for the second one that i did, i need to multiply it by √3! whoa i was right all this time, smh thank you very much!

OpenStudy (ikram002p):

good :) ur wlc

OpenStudy (ikram002p):

it was nice qs